# Extremal Process of Last Progeny Modified Branching Random Walks

### v1

@@ 3.2 BRW in the boundary case
$\operatorname{\mathbf{E}}\left((\widehat{m}_{n}-\widehat{T}_{n})^{\gamma}\mathbbm{1}_{\left\{\widehat{m}_{n}-\widehat{T}_{n}\in[a_{n}-\varepsilon^{-1},\varepsilon n^{1/2}],-\widehat{T}_{j}\geq-A,j\leq n\right\}}\right)\\ \leq\frac{2Ch^{\gamma}(A+1)^{2}(C^{\prime}+h)}{n^{3/2}}\left((\log n)^{\gamma+2}+\sum_{k={\left\lceil\log n\right\rceil}}^{{\left\lceil\varepsilon n^{1/2}/h\right\rceil}}k^{\gamma+1}\right)\\ \leq\frac{K_{\gamma}}{n^{3/2}}(\log n)^{\gamma+2}+\frac{K^{\prime}_{\gamma}}{n^{3/2}}\varepsilon^{\gamma+2}n^{\gamma/2+1},$
@@
[…]
@@
~$\widetilde{\mathcal{E}}_{n}^{(\varepsilon)}=\sum_{|u|=n}\delta_{S_{u}+Y_{u}}\mathbbm{1}_{\left\{S_{u}\leq m_{n}-\varepsilon n^{1/2}\right\}}.$
Using Lemma 3.1, together with the inequality $|e^{-a}-e^{-b}|\leq|a-b|\wedge 1$, we observe that for all $\varphi\in\mathcal{T}$,
$\limsup_{\varepsilon\to 0}\limsup_{n\to\infty}\left|\operatorname{\mathbf{E}}\left(e^{-{\langle\tau_{-m_{n}}\mathcal{E}_{n},\varphi\rangle}}\right)-\operatorname{\mathbf{E}}\left(e^{-{\langle\tau_{-m_{n}}\widetilde{\mathcal{E}}_{n}^{(\varepsilon)},\varphi\rangle}}\right)\right|\\ \leq\lim_{\varepsilon\to 0}\limsup_{n\to\infty}\operatorname{\mathbf{E}}\left(\left(\sum_{|u|=n}\varphi(S_{u}+Y_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}\geq m_{n}-\varepsilon n^{1/2}\right\}}\right)\wedge 1\right)=0.$ (3.8)
It is therefore enough to study the asymptotic behaviour of $\operatorname{\mathbf{E}}\left(e^{-{\langle\tau_{-m_{n}}\widetilde{\mathcal{E}}_{n}^{(\varepsilon)},\varphi\rangle}}\right)$ to identify the limiting distribution of $\tau_{-m_{n}}\mathcal{E}_{n}$.
@@
$\operatorname{\mathbf{E}}\left(\exp\left(-{\langle\tau_{-m_{n}}\widetilde{\mathcal{E}}^{(\varepsilon)}_{n},\varphi\rangle}\right)\right)\\ =\operatorname{\mathbf{E}}\left(\exp\left(-\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}\leq m_{n}-\varepsilon n^{1/2}\right\}}\right)\right).$
We study the asymptotic behaviour of $\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}\leq m_{n}-\varepsilon n^{1/2}\right\}}$ as $n\to\infty$ then $\varepsilon\to 0$. By Lemma 2.3, we get that for all $\delta>0$, for all large enough $n$,
$(1-\delta)c_{\varphi}(\theta)n^{1/2}\sum_{|u|=n}\frac{L(m_{n}-S_{u})}{L(n^{1/2})}e^{\theta S_{u}-n\kappa(\theta)}\mathbbm{1}_{\left\{m_{n}-S_{u}\geq\varepsilon n^{1/2}\right\}}$
@@
$\qquad\qquad\leq(1+\delta)c_{\varphi}(\theta)n^{1/2}\sum_{|u|=n}\frac{L(m_{n}-S_{u})}{L(n^{1/2})}e^{\theta S_{u}-n\kappa(\theta)}\mathbbm{1}_{\left\{m_{n}-S_{u}\geq\varepsilon n^{1/2}\right\}}.$
Recall that $L$ is regularly varying with index $\alpha<0$. We use again [9, Theorem 1.5.2], yielding
@@
$(1-\delta)c_{\varphi}(\theta)n^{1/2}\sum_{|u|=n}\underline{h}_{\varepsilon,\delta}((m_{n}-S_{u})/n^{1/2})e^{\theta S_{u}-n\kappa(\theta)}$
@@
$\qquad\qquad\leq(1+\delta)c_{\varphi}(\theta)n^{1/2}\sum_{|u|=n}\overline{h}_{\varepsilon,\delta}((m_{n}-S_{u})/n^{1/2})e^{\theta S_{u}-n\kappa(\theta)},$
where $\underline{h}_{\varepsilon,\delta}$ and $\overline{h}_{\varepsilon,\delta}$ are continuous functions satisfying for all $x\in\mathbb{R}$
@@
As a result, using a combination of [15, Theorem 1.2] and [4, Theorem 1.1], then letting $\delta\downarrow 0$ we conclude that
@@
$\quad=\sqrt{\frac{2}{\pi\kappa^{\prime\prime}(\theta)}}c_{\varphi}(\theta)Z_{\infty}\operatorname{\mathbf{E}}\left(\left(\sqrt{\kappa^{\prime\prime}(\theta)}R_{1}\right)^{\alpha}\mathbbm{1}_{\left\{\sqrt{\kappa^{\prime\prime}(\theta)}R_{1}\geq\varepsilon\right\}}\right)\quad\text{ in probability,}$,
where $(R_{t},t\in[0,1])$ is a Brownian meander. Since $R_{1}$ has Rayleigh distribution, we observe that $\operatorname{\mathbf{E}}\left(\left(R_{1}\right)^{\alpha}\right)<\infty$. Now, letting $\varepsilon\downarrow 0$, we obtain by monotonicity that
$\lim_{\varepsilon\to 0}\lim_{n\to\infty}\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{m_{n}-S_{u}\geq\varepsilon n^{1/2}\right\}}\\ =\sqrt{\frac{2}{\pi\kappa^{\prime\prime}(\theta)}}c_{\varphi}(\theta)Z_{\infty}\operatorname{\mathbf{E}}\left(\left(\sqrt{\kappa^{\prime\prime}(\theta)}R_{1}\right)^{\alpha}\right)\quad\text{ in probability.}$
@@
$c_{2}=\sqrt{\frac{2}{\pi\kappa^{\prime\prime}(\theta)}}\operatorname{\mathbf{E}}\left(\left(\sqrt{\kappa^{\prime\prime}(\theta)}R_{1}\right)^{\alpha}\right)=\sqrt{\frac{2}{\pi\kappa^{\prime\prime}(\theta)}}\left(2\kappa^{\prime\prime}(\theta)\right)^{\frac{\alpha}{2}}\Gamma\left(\frac{\alpha}{2}+1\right).$
We now conclude, by equation (3.8) that
$\operatorname{\mathbf{E}}(e^{-c_{2}c_{\varphi}(\theta)Z_{\infty}})$ $=\lim_{\varepsilon\to 0}\lim_{n\to\infty}\operatorname{\mathbf{E}}(e^{-{\langle\tau_{-m_{n}}\widetilde{\mathcal{E}}^{(\varepsilon)}_{n},\varphi\rangle}})$
$=\lim_{n\to\infty}\operatorname{\mathbf{E}}(e^{-{\langle\tau_{-m_{n}}\mathcal{E}_{n},\varphi\rangle}}).$
Identifying the Laplace transform of the PPP($c_{2}\theta Z_{\infty}e^{-\theta x}\,\mathrm{d}x$), the proof of Theorem 1.4 is now complete. ∎
3.3 BRW above the boundary case
We prove in this section Theorem 1.7 as a consequence of the convergence of the extremal process of the BRW observed by Madaule [16]. Recall that by [16, Theorem 1.1], for all $\varphi\in\mathcal{T}$, we have
$\lim_{n\to\infty}\operatorname{\mathbf{E}}\left(e^{-{\langle\mathcal{Z}_{n},\varphi\rangle}}\right)=\operatorname{\mathbf{E}}\left(e^{-{\langle\mathcal{Z}_{\infty},\varphi\rangle}}\right),$
where $\mathcal{Z}_{\infty}$ is the limiting extremal process of the BRW, a decorated, randomly shifted Poisson point process with exponential intensity. Moreover, [16, Proposition 1.3] shows that for all $\theta>\theta_{0}$, we have
$\lim_{n\to\infty}\operatorname{\mathbf{E}}\left(\exp\left(-\sum_{|u|=n}e^{\theta(S_{u}-m_{n})}\right)\right)=\operatorname{\mathbf{E}}\left(e^{-{\langle\mathcal{Z}_{\infty},\mathrm{e}_{\theta}\rangle}}\right),$
where $\mathrm{e}_{\theta}:x\mapsto e^{\theta x}$. Using these two results and assuming (1.9), we show that for all $\varphi\in\mathcal{T}$,
$\lim_{n\to\infty}\operatorname{\mathbf{E}}\left(e^{-{\langle\mathcal{Z}_{n},g_{\varphi}\rangle}}\right)=\operatorname{\mathbf{E}}\left(e^{-{\langle\mathcal{Z}_{\infty},g_{\varphi}\rangle}}\right),$
@@ 3.3 BRW above the boundary case
$\left|\operatorname{\mathbf{E}}\left(e^{-{\langle\tau_{-m_{n}}\mathcal{E}_{n},\varphi\rangle}}\right)-\operatorname{\mathbf{E}}\left(e^{-{\langle\mathcal{Z}_{\infty},g_{\varphi}\rangle}}\right)\right|$
[…]
@@
$I_{1}(n,K)$ $=\left|\operatorname{\mathbf{E}}\left(e^{-\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})}\right)-\operatorname{\mathbf{E}}\left(e^{-\sum_{|u|=n}\chi_{K}g_{\varphi}(S_{u}-m_{n})}\right)\right|$
@@
$I_{2}(n,K)$ $=\left|\operatorname{\mathbf{E}}\left(e^{-\sum_{|u|=n}\chi_{K}g_{\varphi}(S_{u}-m_{n})}\right)-\operatorname{\mathbf{E}}\left(e^{-{\langle\mathcal{Z}_{\infty},\chi_{K}g_{\varphi}\rangle}}\right)\right|$
@@
$I_{3}(K)$ $=\left|\operatorname{\mathbf{E}}\left(e^{-{\langle\mathcal{Z}_{\infty},\chi_{K}g_{\varphi}\rangle}}\right)-\operatorname{\mathbf{E}}\left(e^{-{\langle\mathcal{Z}_{\infty},g_{\varphi}\rangle}}\right)\right|.$
@@
Observe that $\chi_{K}g_{\varphi}\in\mathcal{T}$. Therefore, by [16, Theorem 1.1], we know that for all $K>0$, $\lim_{n\to\infty}I_{2}(n,K)=0$. Additionally, $\lim_{K\to\infty}I_{3}(K)=0$ by Lebesgue’s dominated convergence theorem. To complete the proof of Theorem 1.7, we now bound $I_{1}(n,K)$ in $n$.
Let $\theta_{1}\in(\theta_{0},\theta)$, using the inequality $|e^{-a}-e^{-b}|\leq|a-b|\wedge 1$, we have
$I_{1}(n,K)$ $\leq\operatorname{\mathbf{E}}\left(\left(\sum_{|u|=n}(1-\chi_{K})g_{\varphi}(S_{u}-m_{n})\right)\wedge 1\right)$
$\leq\operatorname{\mathbf{E}}\left(\left(\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{|S_{u}-m_{n}|\geq K\right\}}\right)\wedge 1\right)$
$\leq\operatorname{\mathbf{E}}\left(\left(\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}-m_{n}\leq-K\right\}}\right)\wedge 1\right)$
@@
$\qquad\qquad+\operatorname{\mathbf{E}}\left(\left(\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}-m_{n}\geq K\right\}}\right)\wedge 1\right),$
by sub-additivity of $x\mapsto x\wedge 1$. We first observe that
$\operatorname{\mathbf{E}}\left(\left(\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}-m_{n}\geq K\right\}}\right)\wedge 1\right)\leq\operatorname{\mathbf{P}}(M_{n}-m_{n}\geq K),$
@@
which converges to 00 uniformly in $n$ as $K\to\infty$, by tightness of $(M_{n}-m_{n})$. Moreover, using Lemma 2.4, we have
$\operatorname{\mathbf{E}}\left(\left(\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}-m_{n}\leq-K\right\}}\right)\wedge 1\right)$
[…]
$\qquad\quad+\operatorname{\mathbf{P}}\left(\sum_{|u|=n}e^{\theta_{1}(S_{u}-m_{n})}\geq K\right),$
for an arbitrary $\theta_{1}\in(\theta_{0},\theta)$. As a result
@@
$\operatorname{\mathbf{E}}\left(\left(\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}-m_{n}\leq-K\right\}}\right)\wedge 1\right)\\ \leq C^{\prime}Ke^{-K(\theta-\theta_{1})}+\operatorname{\mathbf{P}}\left(\sum_{|u|=n}e^{\theta_{1}(S_{u}-m_{n})}\geq K\right).$
@@
Using [16, Proposition 1.3], we observe that $\sum_{|u|=n}e^{\theta_{1}(S_{u}-m_{n})}$ converges in distribution, thus is tight. As a result, we conclude that
@@
We conclude that
$\lim_{n\to\infty}\operatorname{\mathbf{E}}\left(e^{-{\langle\tau_{-m_{n}}\mathcal{E}_{n},\varphi\rangle}}\right)$ $=\operatorname{\mathbf{E}}\left(e^{-{\langle\mathcal{Z}_{\infty},g_{\varphi}\rangle}}\right)=\operatorname{\mathbf{E}}\left(\exp\left(-\sum_{i\in\mathbb{N}}\varphi(z_{i}+Y_{i})\right)\right)$

### v2

@@ 3.2. BRW in the boundary case
$\operatorname{\mathbf{E}}\left[(\widehat{m}_{n}-\widehat{T}_{n})^{\gamma}\mathbbm{1}_{\left\{\widehat{m}_{n}-\widehat{T}_{n}\in[a_{n}-\varepsilon^{-1},\varepsilon n^{1/2}],-\widehat{T}_{j}\geq-A,j\leq n\right\}}\right]$
$\qquad\leq\frac{2Ch^{\gamma}(A+1)^{2}(C^{\prime}+h)}{n^{3/2}}\left((\log n)^{\gamma+2}+\sum_{k={\left\lceil\log n\right\rceil}}^{{\left\lceil\varepsilon n^{1/2}/h\right\rceil}}k^{\gamma+1}\right)$
$\qquad\leq\frac{K_{\gamma}}{n^{3/2}}(\log n)^{\gamma+2}+\frac{K^{\prime}_{\gamma}}{n^{3/2}}\varepsilon^{\gamma+2}n^{\gamma/2+1},$
@@
$\qquad\leq 2\mathrm{e}^{\theta B}\limsup_{n\to\infty}n^{1/2}\operatorname{\mathbf{E}}\left[\frac{L(\widehat{m}_{n}-\widehat{T}_{n})}{L(n^{1/2})}\mathbbm{1}_{\left\{\widehat{m}_{n}-\widehat{T}_{n}\in[a_{n}-\varepsilon^{-1},\varepsilon n^{1/2}],-\widehat{T}_{j}\geq-A,j\leq n\right\}}\right]$
$\qquad\leq 2\mathrm{e}^{\theta B}K^{\prime}_{\alpha}\varepsilon^{\alpha+2}+2\mathrm{e}^{\theta B}\delta K^{\prime}_{\alpha-\rho}\varepsilon^{\alpha-\rho+2}.$
@@
$\widetilde{\mathcal{E}}_{n}^{(\varepsilon)}=\sum_{|u|=n}\delta_{S_{u}+Y_{u}}\mathbbm{1}_{\left\{S_{u}\leq m_{n}-\varepsilon n^{1/2}\right\}}.$
Using Lemma 3.1, together with the inequality $|\mathrm{e}^{-a}-\mathrm{e}^{-b}|\leq|a-b|\wedge 1$, we observe that for all $\varphi\in\mathcal{T}$,
$\limsup_{\varepsilon\downarrow 0}\limsup_{n\to\infty}\left|\operatorname{\mathbf{E}}\left[\mathrm{e}^{-{\langle\tau_{-m_{n}}\mathcal{E}_{n},\varphi\rangle}}\right]-\operatorname{\mathbf{E}}\left[\mathrm{e}^{-{\langle\tau_{-m_{n}}\widetilde{\mathcal{E}}_{n}^{(\varepsilon)},\varphi\rangle}}\right]\right|$ (3.8)
$\qquad\leq\lim_{\varepsilon\downarrow 0}\limsup_{n\to\infty}\operatorname{\mathbf{E}}\left[\left(\sum_{|u|=n}\varphi(S_{u}+Y_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}\geq m_{n}-\varepsilon n^{1/2}\right\}}\right)\wedge 1\right]=0.$ (3.9)
It is therefore enough to study the asymptotic behaviour of $\operatorname{\mathbf{E}}\left[\mathrm{e}^{-{\langle\tau_{-m_{n}}\widetilde{\mathcal{E}}_{n}^{(\varepsilon)},\varphi\rangle}}\right]$ to identify the limiting distribution of $\tau_{-m_{n}}\mathcal{E}_{n}$.
@@
$\operatorname{\mathbf{E}}\left[\exp\left(-{\langle\tau_{-m_{n}}\widetilde{\mathcal{E}}^{(\varepsilon)}_{n},\varphi\rangle}\right)\right]=\operatorname{\mathbf{E}}\left[\exp\left(-\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}\leq m_{n}-\varepsilon n^{1/2}\right\}}\right)\right].$
We study the asymptotic behaviour of $\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}\leq m_{n}-\varepsilon n^{1/2}\right\}}$ as $n\to\infty$ then $\varepsilon\downarrow 0$. By Lemma 2.2, we get that for all $\delta>0$, for all large enough $n$,
$(1-\delta)c_{\varphi}(\theta)n^{1/2}\sum_{|u|=n}\frac{L(m_{n}-S_{u})}{L(n^{1/2})}\mathrm{e}^{\theta S_{u}-n\kappa(\theta)}\mathbbm{1}_{\left\{m_{n}-S_{u}\geq\varepsilon n^{1/2}\right\}}$
@@
$\qquad\qquad\leq(1+\delta)c_{\varphi}(\theta)n^{1/2}\sum_{|u|=n}\frac{L(m_{n}-S_{u})}{L(n^{1/2})}\mathrm{e}^{\theta S_{u}-n\kappa(\theta)}\mathbbm{1}_{\left\{m_{n}-S_{u}\geq\varepsilon n^{1/2}\right\}}.$
Recall that $L$ is regularly varying with index $\alpha<0$. We use again the Uniform Convergence Theorem (Bingham et al., 1987, Theorem 1.5.2), yielding
@@
$(1-\delta)c_{\varphi}(\theta)n^{1/2}\sum_{|u|=n}\underline{h}_{\varepsilon,\delta}((m_{n}-S_{u})/n^{1/2})\mathrm{e}^{\theta S_{u}-n\kappa(\theta)}$
@@
$\qquad\qquad\leq(1+\delta)c_{\varphi}(\theta)n^{1/2}\sum_{|u|=n}\overline{h}_{\varepsilon,\delta}((m_{n}-S_{u})/n^{1/2})\mathrm{e}^{\theta S_{u}-n\kappa(\theta)},$
where $\underline{h}_{\varepsilon,\delta}$ and $\overline{h}_{\varepsilon,\delta}$ are continuous functions such that, for all $x\in\mathbb{R}$,
@@
As a result, using a combination of (Madaule, 2016, Theorem 1.2) and (Aïdékon and Shi, 2014, Theorem 1.1), then letting $\delta\downarrow 0$ we conclude that
@@
$\quad=\sqrt{\frac{2}{\pi\kappa^{\prime\prime}(\theta)}}c_{\varphi}(\theta)Z_{\infty}\operatorname{\mathbf{E}}\left[\left(R_{1}\sqrt{\kappa^{\prime\prime}(\theta)}\right)^{\alpha}\mathbbm{1}_{\left\{R_{1}\sqrt{\kappa^{\prime\prime}(\theta)}\geq\varepsilon\right\}}\right]\quad\text{ in probability,}$
where $(R_{t},t\in[0,1])$ is a Brownian meander. Since $R_{1}$ has Rayleigh distribution, we observe that $\operatorname{\mathbf{E}}\left[\left(R_{1}\right)^{\alpha}\right]<\infty$. Now, letting $\varepsilon\downarrow 0$, we obtain by monotonicity that
$\lim_{\varepsilon\downarrow 0}\lim_{n\to\infty}\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{m_{n}-S_{u}\geq\varepsilon n^{1/2}\right\}}$
$\qquad=\sqrt{\frac{2}{\pi\kappa^{\prime\prime}(\theta)}}c_{\varphi}(\theta)Z_{\infty}\operatorname{\mathbf{E}}\left[\left(R_{1}\sqrt{\kappa^{\prime\prime}(\theta)}\right)^{\alpha}\right]\quad\text{ in probability.}$
@@
$c_{2}=\sqrt{\frac{2}{\pi\kappa^{\prime\prime}(\theta)}}\operatorname{\mathbf{E}}\left[\left(R_{1}\sqrt{\kappa^{\prime\prime}(\theta)}\right)^{\alpha}\right]=\sqrt{\frac{2}{\pi\kappa^{\prime\prime}(\theta)}}\left(2\kappa^{\prime\prime}(\theta)\right)^{\frac{\alpha}{2}}\Gamma\left(\frac{\alpha}{2}+1\right).$
We now conclude, by equation (3.8), that
$\operatorname{\mathbf{E}}\left[\mathrm{e}^{-c_{2}c_{\varphi}(\theta)Z_{\infty}}\right]=\lim_{\varepsilon\downarrow 0}\lim_{n\to\infty}\operatorname{\mathbf{E}}\left[\mathrm{e}^{-{\langle\tau_{-m_{n}}\widetilde{\mathcal{E}}^{(\varepsilon)}_{n},\varphi\rangle}}\right]=\lim_{n\to\infty}\operatorname{\mathbf{E}}\left[\mathrm{e}^{-{\langle\tau_{-m_{n}}\mathcal{E}_{n},\varphi\rangle}}\right].$
Identifying the Laplace transform of the PPP($c_{2}\theta Z_{\infty}\mathrm{e}^{-\theta x}\,\mathrm{d}x$), the proof of Theorem 1.4 is now complete. ∎
3.3. BRW above the boundary case
We prove in this section Theorem 1.7 as a consequence of the convergence of the extremal process of the BRW observed by Madaule (2017). Recall that by (Madaule, 2017, Theorem 1.1), for all $\varphi\in\mathcal{T}$, we have
$\lim_{n\to\infty}\operatorname{\mathbf{E}}\left[\mathrm{e}^{-{\langle\mathcal{Z}_{n},\varphi\rangle}}\right]=\operatorname{\mathbf{E}}\left[\mathrm{e}^{-{\langle\mathcal{Z}_{\infty},\varphi\rangle}}\right],$
where $\mathcal{Z}_{\infty}$ is the limiting extremal process of the BRW, a decorated, randomly shifted Poisson point process with exponential intensity. Moreover, (Madaule, 2017, Proposition 2.1) shows that $\sum_{|u|=n}\mathrm{e}^{\theta(S_{u}-m_{n})}$ is tight for all $\theta>\theta_{0}$. Using these two results and assuming (1.11), we show that for all $\varphi\in\mathcal{T}$,
$\lim_{n\to\infty}\operatorname{\mathbf{E}}\left[\mathrm{e}^{-{\langle\mathcal{Z}_{n},g_{\varphi}\rangle}}\right]=\operatorname{\mathbf{E}}\left[\mathrm{e}^{-{\langle\mathcal{Z}_{\infty},g_{\varphi}\rangle}}\right],$
@@ 3.3. BRW above the boundary case
$\left|\operatorname{\mathbf{E}}\left[\mathrm{e}^{-{\langle\tau_{-m_{n}}\mathcal{E}_{n},\varphi\rangle}}\right]-\operatorname{\mathbf{E}}\left[\mathrm{e}^{-{\langle\mathcal{Z}_{\infty},g_{\varphi}\rangle}}\right]\right|$
$\qquad=\left|\operatorname{\mathbf{E}}\left[\mathrm{e}^{-\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})}\right]-\operatorname{\mathbf{E}}\left[\mathrm{e}^{-{\langle\mathcal{Z}_{\infty},g_{\varphi}\rangle}}\right]\right|\leq I_{1}(n,K)+I_{2}(n,K)+I_{3}(K),$
@@
$I_{1}(n,K)$ $=\left|\operatorname{\mathbf{E}}\left[\mathrm{e}^{-\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})}\right]-\operatorname{\mathbf{E}}\left[\mathrm{e}^{-\sum_{|u|=n}\chi_{K}g_{\varphi}(S_{u}-m_{n})}\right]\right|$
@@
$I_{2}(n,K)$ $=\left|\operatorname{\mathbf{E}}\left[\mathrm{e}^{-\sum_{|u|=n}\chi_{K}g_{\varphi}(S_{u}-m_{n})}\right]-\operatorname{\mathbf{E}}\left[\mathrm{e}^{-{\langle\mathcal{Z}_{\infty},\chi_{K}g_{\varphi}\rangle}}\right]\right|$
@@
$I_{3}(K)$ $=\left|\operatorname{\mathbf{E}}\left[\mathrm{e}^{-{\langle\mathcal{Z}_{\infty},\chi_{K}g_{\varphi}\rangle}}\right]-\operatorname{\mathbf{E}}\left[\mathrm{e}^{-{\langle\mathcal{Z}_{\infty},g_{\varphi}\rangle}}\right]\right|.$
@@
Observe that $\chi_{K}g_{\varphi}\in\mathcal{T}$. Therefore, by (Madaule, 2017, Theorem 1.1), we know that for all $K>0$, $\lim_{n\to\infty}I_{2}(n,K)=0$. Additionally, $\lim_{K\to\infty}I_{3}(K)=0$ by Lebesgue’s dominated convergence theorem. To complete the proof of Theorem 1.7, we now bound $I_{1}(n,K)$ in $n$.
Let $\theta_{1}\in(\theta_{0},\theta)$, using the inequality $|\mathrm{e}^{-a}-\mathrm{e}^{-b}|\leq|a-b|\wedge 1$ for all $a,b\geq 0$, we have
$I_{1}(n,K)$ $\leq\operatorname{\mathbf{E}}\left[\left(\sum_{|u|=n}(1-\chi_{K})g_{\varphi}(S_{u}-m_{n})\right)\wedge 1\right]$
$\leq\operatorname{\mathbf{E}}\left[\left(\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{|S_{u}-m_{n}|\geq K\right\}}\right)\wedge 1\right]$
$\leq\operatorname{\mathbf{E}}\left[\left(\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}-m_{n}\leq-K\right\}}\right)\wedge 1\right]$
@@
$\qquad\qquad+\operatorname{\mathbf{E}}\left[\left(\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}-m_{n}\geq K\right\}}\right)\wedge 1\right],$
by the sub-additivity $(a+b)\wedge 1\leq(a\wedge 1)+(b\wedge 1)$ for all $a,b\geq 0$. We first observe that
$\operatorname{\mathbf{E}}\left[\left(\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}-m_{n}\geq K\right\}}\right)\wedge 1\right]\leq\operatorname{\mathbf{P}}(M_{n}-m_{n}\geq K),$
@@
which converges to 0 uniformly in $n$ as $K\to\infty$, by tightness of $(M_{n}-m_{n})$. Moreover, using Lemma 2.3, we have
$\operatorname{\mathbf{E}}\left[\left(\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}-m_{n}\leq-K\right\}}\right)\wedge 1\right]$
@@
$\qquad\leq\operatorname{\mathbf{E}}\left[\left(C^{\prime}\sum_{|u|=n}\mathrm{e}^{\theta(S_{u}-m_{n})}\mathbbm{1}_{\left\{S_{u}-m_{n}\leq-K\right\}}\right)\wedge 1\right]$
$\qquad\leq C^{\prime}\operatorname{\mathbf{E}}\left[\mathbbm{1}_{\left\{\sum_{|u|=n}\mathrm{e}^{\theta_{1}(S_{u}-m_{n})}\leq K\right\}}\sum_{|u|=n}\mathrm{e}^{\theta(S_{u}-m_{n})}\mathbbm{1}_{\left\{S_{u}-m_{n}\leq-K\right\}}\right]+\operatorname{\mathbf{P}}\left(\sum_{|u|=n}\mathrm{e}^{\theta_{1}(S_{u}-m_{n})}\geq K\right),$
for an arbitrary $\theta_{1}\in(\theta_{0},\theta)$. As a result,
$\operatorname{\mathbf{E}}\left[\left(\sum_{|u|=n}g_{\varphi}(S_{u}-m_{n})\mathbbm{1}_{\left\{S_{u}-m_{n}\leq-K\right\}}\right)\wedge 1\right]\leq C^{\prime}K\mathrm{e}^{-K(\theta-\theta_{1})}+\operatorname{\mathbf{P}}\left(\sum_{|u|=n}\mathrm{e}^{\theta_{1}(S_{u}-m_{n})}\geq K\right).$
@@
Using (Madaule, 2017, Proposition 2.1), we observe that $\sum_{|u|=n}\mathrm{e}^{\theta_{1}(S_{u}-m_{n})}$ is tight, and therefore we conclude that
@@
Finally,
$\lim_{n\to\infty}\operatorname{\mathbf{E}}\left[\mathrm{e}^{-{\langle\tau_{-m_{n}}\mathcal{E}_{n},\varphi\rangle}}\right]$ $=\operatorname{\mathbf{E}}\left[\mathrm{e}^{-{\langle\mathcal{Z}_{\infty},g_{\varphi}\rangle}}\right]=\operatorname{\mathbf{E}}\left[\exp\left(-\sum_{i\in\mathbb{N}}\varphi(z_{i}+Y_{i})\right)\right],$
@@
Acknowledgements
The authors would like to thank the anonymous referees for their careful reading and detailed comments, which have helped to improve the paper.